a lion eats one sheep in 4 hours, a leopard eats one sheep in 5 hours, and a bear eats one sheep in 6 hours. how long would it take the three animals together to eat one sheep?
Answer by user_dude2008(1862)

t(1/4+1/5+1/6)=1

t(15/60+12/60+10/60)=1

t(37/60)=1

t=(1)(60/37)

t=60/37

t=1.622


Ans: take about 1.622 hrs for all three if they work together.